A steel rotating-beam test specimen has an ultimate strength of 1600 MPa. Estimate the life of the

specimen if it is tested at a completely reversed stress amplitude of 900 MPa.


1
Expert's answer
2020-12-07T03:46:00-0500

Ultimate strength of rotating component is ,Sut=1600 MPa, Reversed amplitude =900 MPa= "\sigma_a"

We know that for endurance limit for specimen is

"S_e=0.5 \times \sigma_u=0.5\times 1600=800" MPa

Now we know that for, "\sigma_u>1400 MPa"

we take endurance limit as 700 MPa and strength friction,f=0.77

Now,

we will find value of constant a and b

"a=\frac{(f\times \sigma_u)^2}{S_e}"

"a=a=\frac{(0.77\times 1600)^2}{700}=2169 MPa"

"b=\frac{1}{3}\times log (\frac{f\times \sigma_u}{S_e})"

b=-0.08

Now we know that for Number of cycles or for life span of specimen

"N=(\frac{\sigma_a}{a})^b=(\frac{900}{2169})^{(-0.08)}"


N= 46401